# Rambo: RL-augmented Model-based Whole-body Control for Loco-manipulation

### v2

@@ III-A Single Rigid Body Dynamics for Legged Robots
$\bm{x}\coloneq[\bm{p}\ \bm{v}\ \bm{R}\ {{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{\omega}}],$ (1)
where $\bm{p}\in\mathbb{R}^{3}$ is the position of the body Center of Mass (CoM); $\bm{v}\in\mathbb{R}^{3}$ is the CoM velocity; $\bm{R}\in SO(3)$ is the rotation matrix of the body frame ${}\{\mathcal{B}\}$ expressed in the inertial world frame $\{\mathcal{W}\}$; $\text{det}(\cdot)$ calculates the determinant of a matrix and $\mathbb{I}$ is the 3-by-3 identity matrix. ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{\omega}}\in\mathbb{R}^{3}$ is the angular velocity expressed in the body frame ${}\{\mathcal{B}\}$. Variables without subscript on the left are assumed to be in the inertial frame $\{\mathcal{W}\}$. Additionally, we define another coordinate frame, namely the projected frame $\{\mathcal{P}\}$, which is centered at the projection of CoM onto the ground plane, and whose $x$ and $z$ axes point forward and upward, respectively.
@@
Figure 2: Illustration of the 3D single rigid body model for a quadruped robot. {${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{\mathcal{W}\}}},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{\mathcal{B}\}}},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\{\mathcal{P}\}}}$ denote the inertial world frame, the body frame, and the projected frame, respectively. $\bm{r}_{i}$ is the contact position of limb $i$ relative to the CoM. $\bm{u}_{i}$ is the external reaction force at the contact position.
@@
where $\bm{F}$ and $\bm{\tau}$ are the total force and torque applied at the CoM; $\bm{r}_{i}=\bm{p}_{i}-\bm{p}$ is the contact positions relative to CoM; the []×$[\cdot]_{\times}:\mathbb{R}^{3}\rightarrow\mathfrak{so}(3)$ operator converts the element from $\mathbb{R}^{3}$ to skew-symmetric matrices as $[\bm{a}]_{\times}\bm{b}=\bm{a}\times\bm{b}$ for all $\bm{a},\bm{b}\in\mathbb{R}^{3}$. The inverse of $[\cdot]_{\times}$ is the vee map $[\cdot]^{\vee}:\mathfrak{so}(3)\rightarrow\mathbb{R}^{3}$
@@
$\dot{\bm{x}}=\begin{bmatrix}\dot{\bm{p}}\\ \dot{\bm{v}}\\ \dot{\bm{R}}\\ {{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\dot{\bm{\omega}}}\end{bmatrix}=\begin{bmatrix}\bm{v}\\ \frac{1}{M}\bm{F}+\bm{g}\\ \bm{R}[{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}\bm{\omega}]_{\times}\\ {{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{I}}^{-1}(\bm{R}^{\top}\bm{\tau}-[{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{\omega}}]_{\times}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{I}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{\omega}})\end{bmatrix},$ (3)
where $M$ is the mass of the rigid body; $\bm{g}\in\mathbb{R}^{3}$ is the gravitational acceleration vector; ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{I}}$ is the fixed moment of inertia in the body frame ${}\{\mathcal{B}\}$.
@@
$\begin{bmatrix}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\dot{\bm{v}}}\\ {{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\dot{\bm{\omega}}}\end{bmatrix}=\begin{bmatrix}\frac{1}{M}\bm{R}^{\top}\bm{F}+\bm{R}^{\top}\bm{g}\\ {{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{I}}^{-1}\bm{R}^{\top}\bm{\tau}\end{bmatrix}=\sum_{i=1}^{N}\begin{bmatrix}\frac{1}{M}\\ {{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{I}}^{-1}[{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{r}_{i}}]_{\times}\end{bmatrix}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{B}}}}}{\bm{u}_{i}}+\hat{\bm{g}},$ (4)
@@ III-B Foot Trajectory Generation Strategy
$({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\bm{p}_{\text{land}}})_{i}=({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\bm{p}_{\text{hip}}})_{i}+\frac{T_{\text{stance}}}{2}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\bm{v}_{\text{hip}}})_{i},$ (5)
@@ IV-A Motion Reference Generation
For each control time step, Rambo starts with querying a motion reference for both the base and joint $({\hat{\bm{q}}},{\hat{\dot{\bm{q}}}})$, where ${\hat{\bm{q}}},{\hat{\dot{\bm{q}}}}$ are the desired generalized coordinates and velocities respectively; ${\hat{\bm{q}}}=[{\hat{\bm{p}}}\ {\hat{\bm{R}}}\ \hat{\bm{q}}_{j}]$ are the desired base position, orientation and joint positions; ${\hat{\dot{\bm{q}}}}=[{\hat{\bm{v}}}\ \hat{\bm{\omega}}\ \hat{\dot{\bm{q}}}_{j}]$ are the desired base linear and angular velocities, joint velocities. We use the subscript $j$ to denote the joint dimensions.
To build a framework that allows the user to interactively control the robot in a versatile manner across diverse tasks, we implement a reference generation process from the user command, shown in Figure 4. The input from the user includes the desired base velocity $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{v}_{x},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{v}_{y},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\omega}_{z})$, which consists of the desired forward, side and turning velocities, all expressed in the projected frame $\{\mathcal{P}\}$. Defining the reference in the projected frame $\{\mathcal{P}\}$ allows Rambo to operate with only proprioceptive information, without relying on an accurate estimation of coordinates in the inertial frame.
In addition, categorizing the end-effectors $i\in\mathcal{N}$ into two kinds: locomotion $\mathcal{L}\subset\mathcal{N}$, marked by superscript ${l}$, and manipulation $\mathcal{M}\subset\mathcal{N}$, marked by superscript ${m}$, our kinematic reference generation also accept specification for the desired EE positions ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{p}}^{m}_{i}}\in\mathbb{R}^{3}$ for the limbs responsible for manipulation. By incorporating predefined gait patterns and their contact schedules, ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{p}}^{l}_{i}}$ is generated by interpolating the keyframes $({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\bm{p}_{\text{lift}}},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\bm{p}_{\text{mid}}},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\bm{p}_{\text{land}}})_{i}$ as shown in Section III-B. The reference for each joint $\hat{\bm{q}}_{j}$ is calculated using inverse kinematics (IK)
$\hat{\bm{q}}_{j}=\text{IK}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\bm{p}},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\bm{R}},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{p}}_{1}},...,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{p}}_{N}}).$ (6)
@@
$\begin{split}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{p}}}&={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{{\bm{p}}},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{R}}}={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{{\bm{R}}}\\ {{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{v}}}&=[{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{v}_{x},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{v}_{y},0],{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{\omega}}}=[0,0,{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\omega}_{z}],\hat{\dot{\bm{q}}}_{j}=\bm{0}.\end{split}$ (7)
@@ IV-B Generating Feedforward Torque via WBC
We firstly calculate the base linear and angular acceleration target ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\bm{a}}=({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\bm{a}}_{l},{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\bm{a}}_{a})\in\mathbb{R}^{6}$ using a proportional-derivative controller
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\bm{a}}_{l}$ $=\kappa_{lp}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{p}}}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{{\bm{p}}})+\kappa_{ld}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{v}}}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{{\bm{v}}})$ (8)
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}\hat{\bm{a}}_{a}$ $=\kappa_{ap}\text{log}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{R}}}^{\top}\cdot{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{{\bm{R}}})^{\vee}+\kappa_{ad}({{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{\hat{\bm{\omega}}}-{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{{}_{\mathcal{P}}}}}{{\bm{\omega}}}),$ (9)

### v4

@@ III-A Single Rigid Body Dynamics for Legged Robots
$\bm{x}\coloneq[\bm{p}\ \bm{v}\ \bm{R}\ {{}_{\mathcal{B}}}{\bm{\omega}}],$ (1)
where $\bm{p}\in\mathbb{R}^{3}$ is the position of the body Center of Mass (CoM); $\bm{v}\in\mathbb{R}^{3}$ is the CoM velocity; $\bm{R}\in SO(3)$ is the rotation matrix of the body frame $\{\mathcal{B}\}$ expressed in the inertial world frame $\{\mathcal{W}\}$; $\text{det}(\cdot)$ calculates the determinant of a matrix and $\mathbb{I}$ is the 3-by-3 identity matrix. ${{}_{\mathcal{B}}}{\bm{\omega}}\in\mathbb{R}^{3}$ is the angular velocity expressed in the body frame $\{\mathcal{B}\}$. Variables without subscript on the left are assumed to be in the inertial frame $\{\mathcal{W}\}$. Additionally, we define another coordinate frame, namely the projected frame $\{\mathcal{P}\}$, which is centered at the projection of CoM onto the ground plane, and whose $x$ and $z$ axes point forward and upward, respectively.
@@
Figure 2: Illustration of the 3D single rigid body model for a quadruped robot. $\{\mathcal{W}\},\{\mathcal{B}\},\{\mathcal{P}\}$ denote the inertial world frame, the body frame, and the projected frame, respectively. $\bm{r}_{i}$ is the contact position of limb $i$ relative to the CoM. $\bm{u}_{i}$ is the external reaction force at the contact position.
@@
where $\bm{F}$ and $\bm{\tau}$ are the total force and torque applied at the CoM; $\bm{r}_{i}=\bm{p}_{i}-\bm{p}$ is the contact positions relative to CoM; the $\cdot]_{\times}:\mathbb{R}^{3}\rightarrow\mathfrak{so}(3)$ operator converts the element from $\mathbb{R}^{3}$ to skew-symmetric matrices as $\bm{a}]_{\times}\bm{b}=\bm{a}\times\bm{b}$ for all $\bm{a},\bm{b}\in\mathbb{R}^{3}$. The inverse of $\cdot$$_{\times}$ is the vee map $\cdot]^{\vee}:\mathfrak{so}(3)\rightarrow\mathbb{R}^{3}$ .
@@
$\dot{\bm{x}}=\begin{bmatrix}\dot{\bm{p}}\\ \dot{\bm{v}}\\ \dot{\bm{R}}\\ {{}_{\mathcal{B}}}{\dot{\bm{\omega}}}\end{bmatrix}=\begin{bmatrix}\bm{v}\\ \frac{1}{M}\bm{F}+\bm{g}\\ \bm{R}[{{}_{\mathcal{B}}}\bm{\omega}]_{\times}\\ {{}_{\mathcal{B}}}{\bm{I}}^{-1}(\bm{R}^{\top}\bm{\tau}-[{{}_{\mathcal{B}}}{\bm{\omega}}]_{\times}{{}_{\mathcal{B}}}{\bm{I}}{{}_{\mathcal{B}}}{\bm{\omega}})\end{bmatrix}$, (3)
where $M$ is the mass of the rigid body; $\bm{g}\in\mathbb{R}^{3}$ is the gravitational acceleration vector; ${{}_{\mathcal{B}}}{\bm{I}}$ is the fixed moment of inertia in the body frame $\{\mathcal{B}\}$.
@@
$\begin{bmatrix}{{}_{\mathcal{B}}}{\dot{\bm{v}}}\\ {{}_{\mathcal{B}}}{\dot{\bm{\omega}}}\end{bmatrix}=\begin{bmatrix}\frac{1}{M}\bm{R}^{\top}\bm{F}+\bm{R}^{\top}\bm{g}\\ {{}_{\mathcal{B}}}{\bm{I}}^{-1}\bm{R}^{\top}\bm{\tau}\end{bmatrix}=\sum_{i=1}^{N}\begin{bmatrix}\frac{1}{M}\\ {{}_{\mathcal{B}}}{\bm{I}}^{-1}[{{}_{\mathcal{B}}}{\bm{r}_{i}}]_{\times}\end{bmatrix}{{}_{\mathcal{B}}}{\bm{u}_{i}}+\hat{\bm{g}}$, (4)
@@ III-B Foot Trajectory Generation Strategy
$({{}_{\mathcal{P}}}{\bm{p}_{\text{land}}})_{i}=({{}_{\mathcal{P}}}{\bm{p}_{\text{hip}}})_{i}+\frac{T_{\text{stance}}}{2}({{}_{\mathcal{P}}}{\bm{v}_{\text{hip}}})_{i},$ (5)
@@ IV-A Motion Reference Generation
[…]
To build a framework that allows the user to interactively control the robot in a versatile manner across diverse tasks, we implement a reference generation process from the user command, shown in Figure 4. The input from the user includes the desired base velocity $({{}_{\mathcal{P}}}\hat{v}_{x},{{}_{\mathcal{P}}}\hat{v}_{y},{{}_{\mathcal{P}}}\hat{\omega}_{z})$, which consists of the desired forward, side and turning velocities, all expressed in the projected frame $\{\mathcal{P}\}$. Defining the reference in the projected frame $\{\mathcal{P}\}$ allows Rambo to operate with only proprioceptive information, without relying on an accurate estimation of coordinates in the inertial frame.
In addition, categorizing the end-effectors $i\in\mathcal{N}$ into two kinds: locomotion $\mathcal{L}\subset\mathcal{N}$, marked by superscript l{l}, and manipulation $\mathcal{M}\subset\mathcal{N}$, marked by superscript m{m}, our kinematic reference generation also accept specification for the desired EE positions ${{}_{\mathcal{P}}}{\hat{\bm{p}}^{m}_{i}}\in\mathbb{R}^{3}$ for the limbs responsible for manipulation. By incorporating predefined gait patterns and their contact schedules, ${{}_{\mathcal{P}}}{\hat{\bm{p}}^{l}_{i}}$ is generated by interpolating the keyframes $({{}_{\mathcal{P}}}{\bm{p}_{\text{lift}}},{{}_{\mathcal{P}}}{\bm{p}_{\text{mid}}},{{}_{\mathcal{P}}}{\bm{p}_{\text{land}}})_{i}$ as shown in Section III-B. The reference for each joint $\hat{\bm{q}}_{j}$ is calculated using inverse kinematics (IK)
$\hat{\bm{q}}_{j}=\text{IK}({{}_{\mathcal{P}}}{\bm{p}},{{}_{\mathcal{P}}}{\bm{R}},{{}_{\mathcal{P}}}{\hat{\bm{p}}_{1}},...,{{}_{\mathcal{P}}}{\hat{\bm{p}}_{N}}).$ (6)
@@
$\begin{split}{{}_{\mathcal{P}}}{\hat{\bm{p}}}&={{}_{\mathcal{P}}}{{\bm{p}}},{{}_{\mathcal{P}}}{\hat{\bm{R}}}={{}_{\mathcal{P}}}{{\bm{R}}}\\ {{}_{\mathcal{P}}}{\hat{\bm{v}}}&=[{{}_{\mathcal{P}}}\hat{v}_{x},{{}_{\mathcal{P}}}\hat{v}_{y},0],{{}_{\mathcal{P}}}{\hat{\bm{\omega}}}=[0,0,{{}_{\mathcal{P}}}\hat{\omega}_{z}],\hat{\dot{\bm{q}}}_{j}=\bm{0}.\end{split}$ (7)
@@ IV-B Generating Feedforward Torque via WBC
We firstly calculate the base linear and angular acceleration target ${{}_{\mathcal{P}}}\hat{\bm{a}}=({{}_{\mathcal{P}}}\hat{\bm{a}}_{l},{{}_{\mathcal{P}}}\hat{\bm{a}}_{a})\in\mathbb{R}^{6}$ using a proportional-derivative controller
${{}_{\mathcal{P}}}\hat{\bm{a}}_{l}$ $=\kappa_{lp}({{}_{\mathcal{P}}}{\hat{\bm{p}}}-{{}_{\mathcal{P}}}{{\bm{p}}})+\kappa_{ld}({{}_{\mathcal{P}}}{\hat{\bm{v}}}-{{}_{\mathcal{P}}}{{\bm{v}}})$ (8)
${{}_{\mathcal{P}}}\hat{\bm{a}}_{a}$ $=\kappa_{ap}\text{log}({{}_{\mathcal{P}}}{\hat{\bm{R}}}^{\top}\cdot{{}_{\mathcal{P}}}{{\bm{R}}})^{\vee}+\kappa_{ad}({{}_{\mathcal{P}}}{\hat{\bm{\omega}}}-{{}_{\mathcal{P}}}{{\bm{\omega}}}),$ (9)